calculate the moment generating function for a random variable which has exponential distribution with parameter gamma.
Answer #1

For an exponential distribution with parameter gamma, therefore the distribution of the random variable here is given as:

X \sim exp(\gamma )

The MGF for X now is obtained here as:

M_x(t) = E[e^{xt}]

M_x(t) = \int_{0}^{\infty } e^{xt}f(x ) \ dx

M_x(t) = \int_{0}^{\infty } e^{xt} \gamma e^{-\gamma x} \ dx

M_x(t) = \int_{0}^{\infty } \gamma e^{-x(\gamma - t)} \ dx

M_x(t) = \left | \frac{ -\gamma e^{-x(\gamma - t)}}{(\gamma - t)} \right |_{0}^{\infty }

M_x(t) =\frac{ \gamma }{(\gamma - t)}

This is the required MGF for the given random variable here.